Let x denote the number of biased die with the following probability mass function p(x) 

X:    1    2   3    4   5     6

P(x):  1/6  1/6  1/5   y   1/5   1/6


a) find the value of y 

b) find the cdf of x


1
Expert's answer
2022-02-07T10:50:05-0500

a)


"\sum_ip(x_i)=1"

"1\/6+1\/6+1\/5+y+1\/5+1\/6=1"

"y=1\/10"

b)


"F(1) =P(X\le 1)=P(1)= 1/6"

"F(2) =P(X\le 2)=P(1)+P(2)= 1/3"

"F(3) =P(X\le 3)=P(1)+P(2)+P(3)= 8/15"

"F(4) =P(X\le4)=P(1)+P(2)+P(3)+P(4)"

"= 19\/30"

"F(5) =P(X\le5)=P(1)+P(2)+P(3)+P(4)"

"+P(5)=5\/6"

"F(6)=P(X\le6)=P(1)+P(2)+P(3)+P(4)"

"+P(X=5)+P(X=6)=1"

The cdf is


"F(x)=\begin{cases}
 0 &x<1 \\
 1/6 &1\le x< 2 \\
 1/3 &2\le x< 3 \\
 8/15 & 3\le x< 4 \\
 19/30 & 4\le x< 5 \\
 5/6 &5\le x< 6 \\
 1 & x\ge 6 \\
\end{cases}"